OpenStudy (anonymous):

What is the missing side length of an irregular quadrilateral with side lengths twenty-five point four centimeters, thirty-three point three centimeters and fifty-two point two centimeters, if the perimeter is equal to the perimeter of a regular hexagon with side length twenty-one centimeters? five point nine centimeters thirteen point two centimeters fifteen point one centimeters thirty-six point one centimeters

OpenStudy (radar):

x + 25.4 + 33.3 + 52.2 = 6(21)

OpenStudy (radar):

x + 110.9 = 126 solve for x.